OpenStudy (anonymous):

Find the equation of the hyperbola having center at the origin. Conjugate axis on the y-axis, eccentricity is square root of 41all over 5, equation of one directrix x= 25 square root of 41 all over 41.

OpenStudy (anonymous):

@RadEn can you pls help me?

OpenStudy (anonymous):

@SolomonZelman can i have your help?

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{(x-h)^2}{a} - \frac{(y-k)^2}{b} } \] eccentricity = (distance from center to focus along transverse axis) ------------------------------------------ a (or semi-transverse axis)

OpenStudy (solomonzelman):

No, Idk, I forgot all this staff, sorry

OpenStudy (anonymous):

yeah, it's okay if that's the case. thanks anyway :)

OpenStudy (anonymous):

pls help @mathstudent55